11.12 IRR AND NPV A company is analyzing two mutually exclusive projects, S and L, with...

11.12

IRR AND NPV

A company is analyzing two mutually exclusive projects, S and L, with the following cash flows:

   0 1 2 3 4
Project S -$1,000 $887.39 $260 $10 $5
Project L -$1,000 $10 $250 $420 $780.08

The company's WACC is 10.5%. What is the IRR of the better project? (Hint: The better project may or may not be the one with the higher IRR.) Round your answer to two decimal places.

____%
Answer #1

S:

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=887.39/1.105+260/1.105^2+10/1.105^3+5/1.105^4

=1026.77

NPV=Present value of inflows-Present value of outflows

=1026.77-1000

=$26.77(Approx).

L:

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=10/1.105+250/1.105^2+420/1.105^3+780.08/1.105^4

=1048.31

NPV=Present value of inflows-Present value of outflows

=1048.31-1000

=$48.31(Approx).

Hence L is better having higher NPV

Let irr be x%
At irr,present value of inflows=present value of outflows.

1000=10/1.0x+250/1.0x^2+420/1.0x^3+780.08/1.0x^4

Hence x=irr=12.1%(Approx).